# MedReadMe: A Systematic Study for Fine-grained Sentence Readability in Medical Domain

### v1

@@ Abstract
Medical texts are notoriously challenging to read. Properly measuring their readability is the first step towards making them more accessible. In this paper, we present a systematic study on fine-grained readability measurements in the medical domain at both sentence-level and span-level. We introduce a new dataset MedReadMe, which consists of manually annotated readability ratings and fine-grained complex span annotation for 4,520 sentences, featuring two novel “Google-Easy” and “Google-Hard” categories. It supports our quantitative analysis, which covers 650 linguistic features and automatic complex word and jargon identification. Enabled by our high-quality annotation, we benchmark and improve several state-of-the-art sentence-level readability metrics for the medical domain specifically, which include unsupervised, supervised, and prompting-based methods using recently developed large language models (LLMs). Informed by our fine-grained complex span annotation, we find that adding a single feature, capturing the number of jargon spans, into existing readability formulas can significantly improve their correlation with human judgments. We will publicly release the dataset and code.
@@ 2.1 Data Collection and Preprocessing
According to the US Census Bureau, [^4] https://nces.ed.gov/programs/digest/d21/ more than 90% of U.S. adults complete high school. To ensure our study reflects the background of a broader audience, our study mainly targets people who have completed high school or are entering college, and our dataset is annotated by college students without medical backgrounds.
@@ 2.2 Sentence-level Readability Annotation
Following prior work Arase et al. (2022), we use the Common European Framework of Reference for Languages (CEFR), which is the most widely used international criteria to define learners’ language proficiency. CEFR assesses language skills on a 6-level scale with a detailed guideline, [^5] https://tinyurl.com/CEFR-Standard/ from beginners to advanced mastery (1: easiest; 6: hardest). As medical texts naturally concentrate on the harder-to-understand side, we introduce the use of “+” and “-” signs to differentiate the nuance in readability, e.g., “3+” and “3-”, in addition to each integer level. They are treated as 3.3 and 2.7 when converting to the numeric scores.
@@
Average number of characters per token 0.524
@@ 3.1 Why Medical Texts are Hard-to-Read?
Figure 3 plots the relationship between readability and both sentence length (left) and the number of jargon spans (right). On the left, we notice that the lines representing “complex” and “simple” sentences begin to diverge as sentence length exceeds 20 tokens, suggesting that factors beyond length affect the readability. In contrast, a stronger overall correlation between the number of jargon spans and readability is observed in the right figure.
@@ 4.1 Evaluating Existing Readability Metrics
The results are presented in Table 7. For supervised methods, we fine-tune language models on our dataset and existing corpora Naous et al. (2023); Arase et al. (2022); Brunato et al. (2018) to predict the sentence readability. We also evaluate the performance of in-context learning by prompting large language models such as GPT-4 and Llama-2 Touvron et al. (2023) using 5-shot. The prompts are constructed following Naous et al. (2023). More details and the full prompt template are in Appendix H. We find that prompt-based methods achieve competitive results, e.g., GPT-4 outperforms the strongest unsupervised metric RSRS, although they still fall behind supervised methods.
@@
Cochrane 0.908 0.549 0.858 0.899 0.870 0.947 0.842 0.850 0.785 0.882
PNAS 0.780 0.574 0.852 0.820 0.791 0.874 0.780 0.824 0.744 0.873
NIHR Series 0.713 0.580 0.824 0.753 0.706 0.885 0.697 0.687 0.634 0.700
eLife 0.538 0.127 0.594 0.715 0.608 0.712 0.812 0.802 0.777 0.861
PLOS Series 0.672 0.309 0.680 0.691 0.635 0.702 0.787 0.843 0.744 0.850
Wiki 0.670 0.429 0.824 0.709 0.607 0.843 0.712 0.619 0.673 0.709
MSD 0.766 0.328 0.784 0.778 0.757 0.867 0.918 0.880 0.863 0.937
Mean ± Std 0.721 $\pm$ 0.115 0.414 $\pm$ 0.17 0.774 $\pm$ 0.1 0.766 $\pm$ 0.073 0.711 $\pm$ 0.101 0.833 $\pm$ 0.092 0.793 $\pm$ 0.076 0.786 $\pm$ 0.096 0.746 $\pm$ 0.075 0.830 $\pm$ 0.090
@@ 4.2 Improving Readability Metrics with Jargon Identification
Length-Controlled Experiment.
5 Fine-grained Complex Span Identification
Data and Models.
@@ 5 Fine-grained Complex Span Identification
Evaluation Metrics.
We consider two variants of F1 measurements: (1) entity-level partial match, indicating the number of jargon, where the type of the predicted entity matches the gold entity and the predicted boundary overlaps with the gold span; [^6] We use the evaluation script released by Tabassum et al. (2020): https://github.com/jeniyat/WNUT_2020_NER/tree/master/code/eval. We also report the exact match performance at entity-level in the Appendix F. (2) token-level match, measuring the number of jargon tokens. For each metric, we conduct evaluations at three levels of granularity: (1) fine-grained level with 7 categories, (2) associated 3 higher-level classes (i.e., medical / general+multisense / abbreviation), and (3) binary judgments between complex or non-complex text spans.
Results.
@@ 7 Conclusion
Due to the reality that major scientific medical discoveries are mostly reported in English, our study primarily focuses on English-language medical texts. Future research could extend to medical resources in other languages. In addition, the readability ratings of a sentence can be impacted by a mixture of factors, including sentence length, grammatical complexity, word difficulty, the annotator’s educational background, the design and quality of annotation guidelines, as well as the target audience. We choose to use the CEFR standards, which is “the most widely used international standard” to access learners’ language proficiency Arase et al. (2022). It has detailed guidelines in 34 languages777 http://tinyurl.com/CEFR-Standard, [^8] http://tinyurl.com/CEFR-34-languages and have been widely used in many prior research (Boyd et al., 2014; Rysová et al., 2016; François et al., 2016; Xia et al., 2016; Tack et al., 2017; Wilkens et al., 2018; Arase et al., 2022; Naous et al., 2023, inter alia).
@@ Acknowledgments
The authors would like to thank Mithun Subhash, Jeongrok Yu, and Vishnu Suresh for their help in data annotation. This research is supported in part by NSF CAREER Award IIS-2144493, NSF Award IIS-2112633, ODNI and IARPA via the HIATUS program (contract 2022-22072200004). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Formulas of Readability Metrics
Here, $y_{p}$ is the predicted distribution from the language model, and $y_{t}$ is the empirical distribution, where 1 for words that appear in the text, and 0 for all others.
@@ Appendix B More Results on the Influence of Each Linguistic Feature
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ Abstract
Medical texts are notoriously challenging to read. Properly measuring their readability is the first step towards making them more accessible. In this paper, we present a systematic study on fine-grained readability measurements in the medical domain at both sentence-level and span-level. We introduce a new dataset MedReadMe, which consists of manually annotated readability ratings and fine-grained complex span annotation for 4,520 sentences, featuring two novel “Google-Easy” and “Google-Hard” categories. It supports our quantitative analysis, which covers 650 linguistic features and automatic complex word and jargon identification. Enabled by our high-quality annotation, we benchmark and improve several state-of-the-art sentence-level readability metrics for the medical domain specifically, which include unsupervised, supervised, and prompting-based methods using recently developed large language models (LLMs). Informed by our fine-grained complex span annotation, we find that adding a single feature, capturing the number of jargon spans, into existing readability formulas can significantly improve their correlation with human judgments. The data is available at tinyurl.com/medreadme-repo.
@@ 2.1 Data Collection and Preprocessing
To ensure our study reflects the background of a broader audience, our study mainly targets people who have completed high school or are entering college, and our dataset is annotated by college students without medical backgrounds using a six-point Likert scale.
@@ 2.2 Sentence-level Readability Annotation
Annotation Guidelines.
Following prior work Arase et al. (2022), we adopt the Common European Framework of Reference for Languages (CEFR) to annotate the sentence readability. CEFR standards were originally created for language learners. Because the scale is essentially a six-point Likert scale, we believe the findings would be mostly generalizable to a broader audience, including native speakers. Another reason for using the CEFR scale is to make our work comparable to the existing work and datasets which were created using the CEFR standards.
@@
CEFR is the most widely used international criteria to define learners’ language proficiency, assessing language skills on a 6-level scale with detailed guidelines, [^4] https://tinyurl.com/CEFR-Standard/ from beginners (A1) to advanced mastery (C2), which are denoted as level 1 (easiest) to level 6 (hardest) in our interface. Following prior work Arase et al. (2022); Naous et al. (2023), a sentence’s readability is determined based on the CEFR level, at which an individual can understand the sentence without assistance. As medical texts naturally concentrate on the harder-to-understand side, we introduce the use of “+” and “-” signs to differentiate the nuance in readability, e.g., “3+” and “3-”, in addition to each integer level. They are treated as 3.3 and 2.7 when converting to the numeric scores.
@@
Number of unique words 0.574
@@ 3.1 Why Medical Texts are Hard-to-Read?
Figure 3 plots the relationship between readability and both the number of jargon spans (right) and sentence length (left), where sentences are split by whitespace. On the left, we notice that the lines representing “complex” and “simple” sentences begin to diverge as sentence length exceeds 20 tokens, suggesting that factors beyond length affect the readability. In contrast, a stronger overall correlation between the number of jargon spans and readability is observed in the right figure.
@@ 4.1 Evaluating Existing Readability Metrics
The results are presented in Table 7. For supervised methods, we fine-tune language models on our dataset and existing corpora Naous et al. (2023); Arase et al. (2022); Brunato et al. (2018) to predict the sentence readability. We also evaluate the performance of in-context learning by prompting large language models such as GPT-4 and Llama-3555More specifically, we used gpt-4-0613 and Llama-3.1-8B-Instruct in the experiments. AI@Meta (2024) using 5-shot. The prompts are constructed following Naous et al. (2023). More details and the full prompt template are in Appendix H. We find that prompt-based methods achieve competitive results, e.g., GPT-4 outperforms the strongest unsupervised metric RSRS, although they still fall behind supervised methods.
@@
Cochrane 0.908 0.665 0.858 0.899 0.870 0.947 0.842 0.850 0.785 0.882
PNAS 0.780 0.528 0.852 0.820 0.791 0.874 0.780 0.824 0.744 0.873
NIHR Series 0.713 0.485 0.824 0.753 0.706 0.885 0.697 0.687 0.634 0.700
eLife 0.538 0.188 0.594 0.715 0.608 0.712 0.812 0.802 0.777 0.861
PLOS Series 0.672 0.520 0.680 0.691 0.635 0.702 0.787 0.843 0.744 0.850
Wiki 0.670 0.447 0.824 0.709 0.607 0.843 0.712 0.619 0.673 0.709
MSD 0.766 0.562 0.784 0.778 0.757 0.867 0.918 0.880 0.863 0.937
Mean ± Std 0.721 $\pm$ 0.115 0.485 $\pm$ 0.148 0.774 $\pm$ 0.1 0.766 $\pm$ 0.073 0.711 $\pm$ 0.101 0.833 $\pm$ 0.092 0.793 $\pm$ 0.076 0.786 $\pm$ 0.096 0.746 $\pm$ 0.075 0.830 $\pm$ 0.090
@@ 4.2 Improving Readability Metrics with Jargon Identification
Length-Controlled Experiment.
5 Fine-grained Complex Span Identification
Data and Models.
Evaluation Metrics.
We consider two variants of F1 measurements: (1) entity-level partial match, indicating the number of jargon, where the type of the predicted entity matches the gold entity and the predicted boundary overlaps with the gold span. We use the evaluation script released by Tabassum et al. (2020).666 https://github.com/jeniyat/WNUT_2020_NER/tree/master/code/eval We also report the exact match performance at entity-level in the Appendix F. (2) token-level match, measuring the number of jargon tokens. For each metric, we conduct evaluations at three levels of granularity: (1) fine-grained level with 7 categories, (2) associated 3 higher-level classes (i.e., medical / general+multisense / abbreviation), and (3) binary judgments between complex or non-complex text spans.
Results.
@@ 7 Conclusion
Due to the reality that major scientific medical discoveries are mostly reported in English, our study primarily focuses on English-language medical texts. Future research could extend to medical resources in other languages. In addition, the focus of our work is to create readability datasets for general purposes following prior work. We did not study or distinguish the fine-grained differences and nuances between native speakers and non-native speakers Yimam et al. (2017).
The readability ratings of a sentence can be impacted by a mixture of factors, including sentence length, grammatical complexity, word difficulty, the annotator’s educational background, the design and quality of annotation guidelines, as well as the target audience. We choose to use the CEFR standards, which is “the most widely used international standard” to access learners’ language proficiency Arase et al. (2022). It has detailed guidelines in 34 languages777 http://tinyurl.com/CEFR-Standard, [^8] http://tinyurl.com/CEFR-34-languages and have been widely used in many prior research (Boyd et al., 2014; Rysová et al., 2016; François et al., 2016; Xia et al., 2016; Tack et al., 2017; Wilkens et al., 2018; Arase et al., 2022; Naous et al., 2023, inter alia).
@@ Acknowledgments
The authors would like to thank Mithun Subhash, Jeongrok Yu, and Vishnu Suresh for their help in data annotation. This research is supported in part by the NSF CAREER Award IIS-2144493, NSF Award IIS-2112633, NIH Award R01LM014600, ODNI and IARPA via the HIATUS program (contract 2022-22072200004). The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF, NIH, ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ Appendix A Formulas of Readability Metrics
Here, $S$ is sentence length, $y_{p}$ is the predicted distribution from the language model, and $y_{t}$ is the empirical distribution, where 1 for words that appear in the text, and 0 for all others.
@@ Appendix B More Results on the Influence of Each Linguistic Feature
LFTK Lee and Lee (2023) corr_adj_var 0.4764 Corrected adjective variation, which is computed as $(\frac{\text{number-of-unique-adjectives}}{\sqrt{2\times\text{number-of-all-adjectives}}})$
@@
LingFeat Lee et al. (2021) SquaAjV_S 0.4088 Squared Adjective Variation-1, which is calculated as the $(\frac{(\text{number-of-unique-adjectives})^{2}}{\text{number-of-total-adjectives}})$.
@@
LFTK Lee and Lee (2023) corr_adp_var 0.4031 Corrected adposition variation, which is computed as $(\frac{\text{number-of-unique-adpositions}}{\sqrt{2\times\text{number-of-all-adpositions}}})$
@@
LFTK Lee and Lee (2023) corr_propn_var 0.3895 Corrected proper noun variation, which is computed as $(\frac{\text{number-of-unique-proper-nouns}}{\sqrt{2\times\text{number-of-all-proper-nouns}}})$